OpenStudy (anonymous):

Solve for x and y 12sinx+5cosx = 2y^2-8y+21

OpenStudy (debbieg):

1 equation in 2 unknowns??

OpenStudy (anonymous):

this is a really good question..tell how you approached it...

OpenStudy (anonymous):

@deltarokker

OpenStudy (anonymous):

i multiplied and divided the equations by \[\sqrt{ a ^{2}+b ^{2}}\]

OpenStudy (anonymous):

i knew the simplified part will have magnitude less than 1 but i was stuck then again

OpenStudy (anonymous):

@RaGhavv

OpenStudy (anonymous):

a=12,b=5?

OpenStudy (anonymous):

@RaGhavv yes

OpenStudy (anonymous):

okay...so on left side you have an expression sumthing lyk13[sin(x+sumthng)],right..? n what do you infer from this..?

OpenStudy (debbieg):

Use reduction formula? then LHS becomes: \[13\sin(x+\tan^{-1} \frac{ 5 }{ 12 })\]

OpenStudy (anonymous):

we know that the value for the trigonometric eqn lies between,\[\left( -\sqrt{a ^{2}+b ^{2}},\sqrt{a ^{2}+b ^{2}} \right)\]

OpenStudy (anonymous):

good...now analyse the RHS part...can you tell something about it..? lyk you said about the left part..?